# Subvarieties of pointed Abelian $\ell$-groups

### v1

@@ 6 Applying Mundici functor
Let $\mathbb{K}$}} be a proper variety of MV-algebras. Denote some ${\boldsymbol{\mathit{A}}}\notin{{\mathbb{K}}}$. Then we have by the lemma above $\Gamma^{-1}({\boldsymbol{\mathit{A}}})\notin\mathbf{HSP}(\Gamma^{-1}[{{\mathbb{K}}}])$. Thus $\mathbf{HSP}(\Gamma^{-1}[{{\mathbb{K}}}])$ is a proper subvariety of $p\mathbb{AL}^{+}$ and thus we have $\mathbf{HSP}(\Gamma^{-1}[{{\mathbb{K}}}])=V_{I,J}$ for some finite sets $J\subseteq I$. Since ${\boldsymbol{\mathit{Z}}}_{i},{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in V_{I,J}$ for $i\in I,j\in J$, by lemma above we obtain $\Gamma({\boldsymbol{\mathit{Z}}}_{i}),\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})\in{{\mathbb{K}}}$ for $i\in I$ and $j\in J$. Thus $\{\Gamma({\boldsymbol{\mathit{Z}}}_{i}),\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})\mid i\in I,j\in J\}\subseteq{{\mathbb{K}}}.$
To show the other inclusion we assume that ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}$. Consequently, $\Gamma^{-1}({\boldsymbol{\mathit{A}}})\in\Gamma^{-1}[{{\mathbb{K}}}]$ and thus $\Gamma^{-1}({\boldsymbol{\mathit{A}}})\in V_{I,J}$ which means ${\boldsymbol{\mathit{A}}}\in\{\Gamma({\boldsymbol{\mathit{Z}}}_{i}),\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})\mid i\in I,j\in J\}$. This proves ${{\mathbb{K}}}=\{\Gamma({\boldsymbol{\mathit{Z}}}_{i}),\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})\mid i\in I,j\in J\}$ for some finite sets $J\subseteq I$. ∎
Similarly, one can show semantical part of Theorem 5.9 using Komori classification and Theorem 3.10. For details one can see discussion in [28]. We prove here something stronger.
Let $\mathbf{S}$ denote any finitely generated dense $\ell$-subgroup of $\mathbf{R}$ such that $\mathbf{S}\cap\mathbf{Q}=\mathbf{Z}$. Recall the result from [14]:
@@
Let $\mathbb{M}$ be a class of MV-chains. Then the following properties are equivalent:
@@
We need to show how Mundici functor behaves with respect to ultrapowers. Generally, it is not true that $\Gamma(\prod_{i\in\omega}{\boldsymbol{\mathit{A}}}_{a}/\mathfrak{U})=\prod_{i\in\omega}\Gamma({\boldsymbol{\mathit{A}}}_{a})/\mathfrak{U}$, where $\mathfrak{U}$ is a non-principal ultrafilter. In fact one can easily show that for strongly pointed $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}$ the $\ell$-group $\prod_{i\in\omega}{\boldsymbol{\mathit{A}}}_{a}/\mathfrak{U}$ is never even strongly pointed. However, we claim that Mundici functor preserves universal subclasses. We will need the following lemma, which we prove later in the end of this section.
@@
Let ${{\mathbb{K}}}\cup\{{\boldsymbol{\mathit{A}}}_{u}\}$ be a class of totally ordered strongly positively pointed Abelian $\ell$-groups. Then we have ${\boldsymbol{\mathit{A}}}_{u}\in\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})$ iff $\Gamma({\boldsymbol{\mathit{A}}}_{u})\in\mathbf{ISP}_{\mathrm{U}}(\Gamma[{{\mathbb{K}}}])$.
Using these tools we can prove the following.
@@
Implication $2$ to 1 is trivial. To prove the other implication assume $\mathbb{M}$ is universal. Therefore, by Lemma 2.3 we have $\mathbb{M}=\mathbf{ISP}_{\mathrm{U}}(\mathbb{M})$. Let us denote $\mathbb{K}$}} the subclass of $M$ consisting of all strongly pointed Abelian $\ell$-groups.
@@
    If ${\boldsymbol{\mathit{Z}}}_{0}\in\mathbb{M}$ then $p\mathbb{AL}_{SI}^{0}\subseteq\mathbb{M}$ by Lemma 2.11 and thus by Theorem 3.10 we have
    $\mathbb{M}=\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})\cup p\mathbb{AL}_{SI}^{0}=\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{0})=$
    $\mathbf{ISP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{Z}}}_{i}\mid i\in I\}\cup\{{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}\mid j\in J,d_{j}\in D_{j}\}\cup\{\mathbf{S}_{k}\mid k\in K\}).$
@@
    If $Z_{0}\notin\mathbb{M}$, we get by Lemma 2.11 that $p\mathbb{AL}_{SI}^{0}\cap\mathbb{M}=\emptyset$ and thus $\mathbb{M}=$
$\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})=\mathbf{ISP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{Z}}}_{i}\mid i\in I\}\cup\{{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}\mid j\in J,d_{j}\in D_{j}\}\cup\{\mathbf{S}_{k}\mid k\in K\}).$
@@
Let us just say that Theorem 6.6 and Corollary 6.7 can be stated also for negatively pointed Abelian $\ell$-groups and thus we can get description of all universal classes of totally ordered pointed Abelian $\ell$-groups. This gives us two corollaries.

### v2

@@ 6 Applying the Mundici functor
Let $\mathcal{V}$ be a proper variety of MV-algebras. Since $\mathcal{V}$ is proper, there is ${\boldsymbol{\mathit{A}}}\notin\mathcal{V}$. Let us consider the variety of positively pointed Abelian $\ell$-groups ${{\mathbb{K}}}=\mathbf{HSP}(\Gamma^{-1}[\mathcal{V}])$. By Lemma 6.2 we have
$\Gamma[{{\mathbb{K}}}]=\Gamma(\mathbf{HSP}(\Gamma^{-1}[\mathcal{V}]))\subseteq\mathbf{HSP}(\Gamma[\Gamma^{-1}[\mathcal{V}]])=\mathbf{HSP}(\mathcal{V})=\mathcal{V}.$
Since ${\boldsymbol{\mathit{A}}}\notin\mathcal{V}$, also $\Gamma(\Gamma^{-1}({\boldsymbol{\mathit{A}}}))\notin\mathcal{V}$ and by Lemma 6.2 we obtain $\Gamma^{-1}({\boldsymbol{\mathit{A}}})\notin{{\mathbb{K}}}$. Therefore, $\mathbb{K}$}} is a proper subvariety of $p\mathbb{AL}^{+}$ and thus by Theorem 5.7 we have $\mathbf{HSP}(\Gamma^{-1}[\mathcal{V}])=\mathcal{V}_{I,J}$ for some finite sets $J\subseteq I$. Since ${\boldsymbol{\mathit{Z}}}_{i},{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathcal{V}_{I,J}$ for $i\in I,j\in J$, by Lemma 6.2 we obtain $\Gamma({\boldsymbol{\mathit{Z}}}_{i}),\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})\in\Gamma[\mathcal{V}_{I,J}]\subseteq\mathcal{V}$ for $i\in I$ and $j\in J$.
Now it remains to check $\mathcal{V}$ is generated by $\Gamma({\boldsymbol{\mathit{Z}}}_{i}),\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$ for every $i\in I,j\in J$. Let ${\boldsymbol{\mathit{B}}}\in\mathcal{V}$. We obtain $\Gamma^{-1}({\boldsymbol{\mathit{B}}})\in\Gamma^{-1}[\mathcal{V}]\subseteq\mathbf{HSP}(\Gamma^{-1}[\mathcal{V}])=\mathcal{V}_{I,J}=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{i},{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\mid i\in I,j\in J).$ By applying $\Gamma$ it follows ${\boldsymbol{\mathit{B}}}\in\mathbf{HSP}(\Gamma({\boldsymbol{\mathit{Z}}}_{i}),\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})\mid i\in I,j\in J).$ This completes the proof. ∎
Here, we used the classification of subvarieties of $p\mathbb{AL}^{+}$ to obtain the classification of all subvarieties of MV-algebras. We should note that we could proceed in the opposite way: using Komori’s classification, Corollary 3.6, and Theorem 3.10, we can derive the semantic part of Theorem 5.9. For further details, see the discussion in [32].
6.1 Gispert’s classification of universal classes of totally ordered MV-algebras
In the rest of this section we focus on quasivarieties and universal classes. Let $\mathbf{S}$ denote any finitely generated dense $\ell$-subgroup of $\mathbf{R}$ such that $\mathbf{S}\cap\mathbf{Q}=\mathbf{Z}$. Recall the result from [16]:
@@ 6.1 Gispert’s classification of universal classes of totally ordered MV-algebras
Let $\mathbb{M}$ be a class of totally ordered MV-algebras. Then the following properties are equivalent:
@@
We need to show how $\Gamma$ behaves with respect to ultraproducts.
@@
Let ${{\mathbb{K}}}\cup\{{\boldsymbol{\mathit{A}}}_{a}\}$ be a class of strongly positively pointed Abelian $\ell$-groups. Then we have ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})$ iff $\Gamma({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}(\Gamma[{{\mathbb{K}}}])$.
Proof.
Throughout this proof, let $I$ be an arbitrary fixed set, $\mathfrak{U}$ be an ultrafilter over $I$ and $\{{\boldsymbol{\mathit{A}}}^{i}_{a_{i}}\mid i\in I\}\subseteq{{\mathbb{K}}}$. First, let us note that $\Gamma(\prod_{i\in I}{\boldsymbol{\mathit{A}}}^{i}_{a_{i}}/\mathfrak{U})=\Gamma(\prod_{i\in I}{\boldsymbol{\mathit{A}}}^{i}_{a_{i}})/\mathfrak{U}=(\prod_{i\in I}\Gamma({\boldsymbol{\mathit{A}}}^{i}_{a_{i}})/\mathfrak{U})$. Consequently, ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})$ implies $\Gamma({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}(\Gamma[{{\mathbb{K}}}])$.
For the other inclusion first assume $\Gamma({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{IS}(\prod_{i\in I}\Gamma({\boldsymbol{\mathit{A}}}^{i}_{a_{i}})/\mathfrak{U})$. Let ${\boldsymbol{\mathit{B}}}_{b}$ denote the strongly pointed convex $\ell$-subgroup of $\prod_{i\in I}{\boldsymbol{\mathit{A}}}^{i}_{a_{i}}/\mathfrak{U}$. We know that we have $\Gamma^{-1}(\Gamma(\prod_{i\in I}{\boldsymbol{\mathit{A}}}^{i}_{a_{i}}/\mathfrak{U}))={\boldsymbol{\mathit{B}}}_{b}$. Since $\Gamma^{-1}$ preserves embeddings, we have $\Gamma^{-1}(\Gamma({\boldsymbol{\mathit{A}}}_{a}))\in\mathbf{IS}(\Gamma^{-1}(\prod_{i\in I}\Gamma({\boldsymbol{\mathit{A}}}^{i}_{a_{i}})/\mathfrak{U}))$. Since ${\boldsymbol{\mathit{A}}}_{a}\cong\Gamma^{-1}(\Gamma({\boldsymbol{\mathit{A}}}_{a}))$ and
$\Gamma^{-1}(\prod_{i\in I}\Gamma({\boldsymbol{\mathit{A}}}^{i}_{a_{i}})/\mathfrak{U})=\Gamma^{-1}(\Gamma(\prod_{i\in I}{\boldsymbol{\mathit{A}}}^{i}_{a_{i}}/\mathfrak{U}))={\boldsymbol{\mathit{B}}}_{b},$
we obtain ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{IS}({\boldsymbol{\mathit{B}}}_{b})$. Since ${\boldsymbol{\mathit{B}}}_{b}\in\mathbf{IS}(\prod_{i\in I}{\boldsymbol{\mathit{A}}}^{i}_{a_{i}}/\mathfrak{U})$, we have ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{IS}(\prod_{i\in I}{\boldsymbol{\mathit{A}}}^{i}_{a_{i}}/\mathfrak{U})$ and thus ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})$. This completes the proof. ∎
@@
The implication $2$ to $1$ is trivial. To prove the other implication assume $\mathbb{M}$ is universal. Therefore, by Theorem 2.3 we have $\mathbb{M}=\mathbf{ISP}_{\mathrm{U}}(\mathbb{M})$. Let us denote by $\mathbb{K}$}} the subclass of $\mathbb{M}$ consisting of all strongly pointed Abelian $\ell$-groups.
@@
    If ${\boldsymbol{\mathit{Z}}}_{0}\in\mathbb{M}$ then $p\mathbb{AL}_{\mathrm{FSI}}^{0}\subseteq\mathbb{M}$ by Lemma 2.13 and thus by Theorem 3.10 we have
    $\mathbb{M}$ $=\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})\cup p\mathbb{AL}_{\mathrm{FSI}}^{0}=\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{0})$
    $=\mathbf{ISP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{Z}}}_{i}\mid i\in I\}\cup\{{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}\mid j\in J,d_{j}\in D_{j}\}\cup\{\mathbf{S}_{k}\mid k\in K\}).$
@@
    If ${\boldsymbol{\mathit{Z}}}_{0}\notin\mathbb{M}$, we obtain by Lemma 2.13 that $p\mathbb{AL}_{\mathrm{FSI}}^{0}\cap\mathbb{M}=\emptyset$ and thus
    $\mathbb{M}$ $=\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})$
    $=\mathbf{ISP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{Z}}}_{i}\mid i\in I\}\cup\{{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}\mid j\in J,d_{j}\in D_{j}\}\cup\{\mathbf{S}_{k}\mid k\in K\}).$
@@
Let us note that analogical statements to Theorem 6.6 and Corollary 6.7 can be proved also for negatively pointed Abelian $\ell$-groups. Therefore, we can generalize the Theorem 6.6 and Corollary 6.7 as follows.
@@
Let $\mathbb{M}$ be a class of totally ordered pointed Abelian $\ell$-groups. Then the following properties are equivalent: